OpenStudy (anonymous):

Can I rewrite 2x=4+z as z=4-2x???

OpenStudy (kira_yamato):

Nope... z = 2x - 4

OpenStudy (anonymous):

no ,because

OpenStudy (anonymous):

hmmm why is the 4 become negative if it is already after the = sign

OpenStudy (anonymous):

*does

OpenStudy (anonymous):

but the 4 is already after the = sign, why does the sign change?

OpenStudy (anonymous):

only the z and 2x swap sides

OpenStudy (anonymous):

the 4 is always left of the = sign though

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

=4 -4= ?????

OpenStudy (anonymous):

What grade are you, Chrissy?

OpenStudy (anonymous):

year 8 uk school

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

how can the -4 still be negative even though its moved

OpenStudy (anonymous):

that goes against everything I have learnt thus far

OpenStudy (anonymous):

Alright but I think understanding these things by moving the variable is not much of intuitive and may cause confusion in new learners. So my attempt would be to proceed in the follow way: You have \[2x=4+z\] and now you want to isolate z so you have to get rid of 4 from the right hand side .. we can do do this my subtracting 4 from the RHS but to preserve the equation you also need to subtract 4 from LHS so . \[2x-4=4+z-4 \Rightarrow 2x-4=z\] Hence your answer.

OpenStudy (anonymous):

@Hollywood_chrissy:Don't move things just read my answer and let me know if you didn't understand :)

OpenStudy (anonymous):

I could kiss you @ foolformaths, @ diyadiya thank you for your patience with me.

OpenStudy (anonymous):

Thanks,and glad to help .. you may try to kiss me but I think my gf is not gonna like it :P :D

OpenStudy (anonymous):

I understand from linear equations that what you do to one side you must do to the other, so foolformaths you made perfect sense.

OpenStudy (anonymous):

@Diya: Moving things gets more messy when dealing with inequalities

OpenStudy (anonymous):

I am sure it is we all have a complex brain to program anything and everything ;) :D